OpenStudy (anonymous):

3x^2-2√(3)xy+y^2+2x+2√(3)y=0 determine the type of conic, solve for y, determine the angle of rotation-use the formula cot2=A-C/B and rewrite the equation of the conic using the equation x=x'cos-y'sin and y=x'sin+y'cos i can't find that! =/

OpenStudy (anonymous):

WOW! You have hard problems... (I am SUCH a big help)

OpenStudy (anonymous):

I don't know how to do this... :,(

OpenStudy (anonymous):

lol, yea it's very complicated =/

OpenStudy (anonymous):

hate math with a passion!

OpenStudy (anonymous):

This seems really advanced. What class is this for?

OpenStudy (anonymous):

it seems as this is a quadratic equation in the form:Ax^2 + Bxy + Cy^ 2 + Dx + Ey + F = 0 now the reason you are asked for the rotation, is because of the Bxy term. Now lets see if we can apply the discriminate B^2-4AC, to determine the type of conic

OpenStudy (anonymous):

So, we have B=-2sqrt(3) and A=3 and C=1, so let do that computation: (-2sqrt(3))^2-4(3)(1) 4(3)-12=12-12=0, thus we have a parabola

OpenStudy (anonymous):

=o=

OpenStudy (anonymous):

Now we have to actually find the angle of rotation, we have at our disposal two equations, although we are told explicitly to use